OpenStudy (anonymous):

How do I solve this system using substitution? In steps please? :) x + 12y = 68 x + 8y - 12

OpenStudy (anonymous):

where the = in second equation?

OpenStudy (anonymous):

Assuming the second equation should be x + 8y = 12...here's what I would do: subtract 8y from each side in eq 2 x = 12 - 8y Then substitute "12 - 8y" into the first equation for x 12 - 8y +12y = 68 Solve for y, then use that result in the original second equation to find x

OpenStudy (anonymous):

Sorry. It's x = 8y - 12

OpenStudy (anonymous):

No problem...then it is already in form to substitute. use "8y - 12" where I said "12 - 8y"

OpenStudy (anonymous):

So for 12 - 8y +12y = 68 , do I add 12 on the other side?

OpenStudy (anonymous):

oh wait do i combine 8y + 12y?

OpenStudy (anonymous):

it is -8y + 12y = 4y so your equation would simplify to 12 + 4y = 68 Then you would subtract 12 from each side to get 4y=56 then to find y just divide both sides by 4 which gives you y=14 Now that you have y you can substitute it in to solve for x. x=8y-12 so x=8(14)-12 which goes to x=112-12 which trivial math gives you x=100 so your solution to your system is (100,14)

OpenStudy (anonymous):

Where did you get 4 ?